# Cluster and Predict Latent Patches for Improved Masked Image Modeling

### v2

@@ 1 Introduction
In this work, we focus on online latent masked image modeling. We isolate it from other stabilizing objectives, and systematically study the design choices it involves. We center the discussion around three aspects of the masked image modeling principle: the target representation, the formulation of the loss, and the architecture used to perform predictions. We show that with the right implementation, a simple masked image modeling objective can lead to features competitive with the current state of the art in SSL. In brief, our method relies on a pair of teacher-student vision transformers trained with self-distillation (fig. 1 left). The training signal for the student comes from the patch representations of the teacher, which is updated through an exponential moving average (EMA). The patch embeddings of the teacher are converted to soft assignments on pseudo-categories using a learned online clustering, while the student receives a partially masked image as input and is trained to predict the clustering assignments of the missing patches. Using this method, we train CAPI, a 300-million parameter visual encoder whose representations approach DINOv2’s performance while significantly outperforming previous masked image models (fig. 1 right).
Figure 1: CAPI Method overview: image patches embedded by a teacher are grouped into clusters. Their assignments are then used as the training signal for the student. The teacher and the student are jointly learned via self-distillation. The loss is purely about predicting the content of missing patches and does not rely on augmentations or a contrastive loss. Evaluation scores: we evaluate frozen representations on ADE20K segmentation with a $k$-nn and linear probe and on ImageNet-1k classification with an attentive probe. We compare to MAE, data2vec 2.0, I-JEPA, and AIM. Compared to other masked image models, CAPI achieves higher performance with fewer FLOP, scaling well with model size, and approaches the scores of DINOv2+reg.
@@ 2 Related Work
Self-supervised representation learning. Self-supervised learning (SSL) is a pre-training paradigm in which a model is optimized to solve a pretext task on unlabeled data, often collected at scale. As the model is not trained to match specific human annotations, the benefit of this type of training is to produce a generalist model, that can be adapted to solve many different downstream tasks. Depending on the application, these tasks can be solved either by fully fine-tuning the mode, or by using the representations extracted with the frozen model. In the SSL literature, some works have focused mainly on full fine-tuning (He et al., 2022; Huang et al., 2023), while others have shown that frozen representations can reach excellent performance on a wide range of tasks, avoiding costly fine-tuning (Oquab et al., 2024) and generalizing to annotation-scarce domains where fine-tuning is not possible (Tolan et al., 2024; Xu et al., 2024). Historically, early work on self-supervised learning focused on hand-crafted pretext tasks such as predicting the rotation of an image (Gidaris et al., 2018), the relative position of patches (Doersch et al., 2015) or the color of a grayscale image (Zhang et al., 2016). Subsequent works pushed the field forward with methods based on clustering (Caron et al., 2018; 2020; 2021) and contrastive learning (Chen et al., 2020b; He et al., 2020). Nowadays, the best self-supervised encoder inherits from these families (Oquab et al., 2024) and complements them with a masked image modeling objective (Zhou et al., 2022). However, training with both global and MIM objectives can prove difficult, as multiple components can interact negatively111 Zhou et al. (2022) showed that using a shared head for the DINO and iBOT losses gave better results at small scales. However, at large scale, Oquab et al. (2024) observed that this caused the iBOT loss to explode late into the training, significantly degrading performance. Untying the two heads prevented the two losses from competing directly, stabilizing the training. . In this work, we study the masked image modeling component in isolation, suggesting improvements to properly stabilize the optimization objective in the absence of a global term.
@@
Latent reconstruction. Concurrently to MAE, Zhou et al. (2022) proposed iBOT. To obtain a more semantic tokenization, iBOT used the online output of the model being trained as the reconstruction target for masked image modeling. This led to good representations, and the method was reused to obtain the current state-of-the-art (Oquab et al., 2024). However, the iBOT objective was very unstable and required an additional DINO (Caron et al., 2021) loss to stabilize the training. The idea of using online representations as targets was then reused in data2vec (Baevski et al., 2022) and I-JEPA (Assran et al., 2023; Bar et al., 2024). I-JEPA in particular proposed reusing the encoder/decoder architecture of MAE and removing the projection head of iBOT, to obtain a more stable objective in the latent space. The improvements in I-JEPA established a new tradeoff between stability and performance, but it was still both sensitive to hyperparameters ($-12$ points on IN-1k when changing the “target scale” from $[0.15,0.2]$ to $[0.125,0.2]$ (Assran et al., 2023))) and weaker than DINOv2 (81.1 on ImageNet-1k, 5.4 below DINOv2 while using a model twice bigger).
Clustering in self-supervised learning. Our approach is also related to methods that use clustering for self-supervised learning. First of this line of work, DeepCluster (Caron et al., 2018) proposed using a simple $k$-means to obtain pseudo-labels. Subsequently, SwAV (Caron et al., 2020) introduced an online clustering to replace the offline $k$-means. DINO (Caron et al., 2021) built on SWaV, and made the clustering be implicitly learned by the MLP projection head of the student. Finally, iBOT (Zhou et al., 2022) proposed to reuse the DINO projection head for masked image modeling, implicitly using a clustering as target. Our approach reuses this idea of using a clustering to create the targets of a masked image modeling objective, but backtracks to the origin of this line of work: instead of using an MLP head that performs an implicit clustering, we use an explicit clustering. By separating the clustering process from the rest of the training, we isolate the two components, making the training more stable and transparent.
@@ 3 Approach
Figure 3: The target representations commonly used in MIM. We focus on the EMA representations.
Overview of training. In short, our main design choices are: first, we reconstruct images in latent space with a teacher-student framework, following iBOT. Then, we formulate our loss using a clustering component, inspired by DeepCluster, SwAV, and DINO, and draw inspiration from the regularization methods they introduce, in particular the Sinkhorn-Knopp (Sinkhorn & Knopp, 1967). Finally, we employ a cross-attention predictor model, separate from the encoder, to perform reconstructions, following crossMAE (Fu et al., 2024).
Our encoder and the predictor are transformers (Dosovitskiy et al., 2021) and, during pre-training, they operate in tandem as the student (fig. 1 left). The teacher is an EMA of the encoder. The typical values for one training iteration using square images of side 224 pixels are:
@@ 3.1 Clustering-based loss formulation
Latent clustering methods such as SwAV and DINO employ a cross-entropy loss between the student and teacher output distributions. These distributions, produced by a linear or MLP head, can be seen as soft cluster memberships, where the centroids correspond to the prototypes. Replicating the DINO objective, iBOT proposes to pass the student predictions through an MLP head and to pass the teacher embeddings through the EMA of this head. But this ignores a specificity of masked image modeling: while in DINO both targets and predictions are [CLS] tokens, in iBOT the targets are patch tokens while the predictions are special [MSK] tokens. This causes a distribution mismatch: the MLP head of the student is trained with [MSK] inputs but is instead applied to regular patch tokens in the teacher. This mismatch would be even stronger in an asymmetric architecture as in MAE or I-JEPA, where the targets and predictions come from two different networks (see fig. 5). Without the stabilizing effect of the DINO loss, the iBOT formulation is unable to bootstrap itself and results in trivial representations (see table 1(c)). Our proposition is simple: we decouple the training of the teacher projection from the student’s head by directly learning an online clustering of the teacher patch tokens. This way, the training remains stable even in the absence of a stabilizing loss.
(a) Direct loss
@@
(c) Clustering
Figure 4: The different loss formulations considered here. We depict in red the flow of the gradient.
@@
(BeiT, iBOT)
(b) Self-att. predictor
@@
Figure 5: The different predictor architectures discussed in the paper. Here, the boxes each represent a transformer. The black lines represent the residual stream for a token.
@@
where $a_{i}^{(k)}$ is the $k$-th coordinate in $a_{i}$. By minimizing the entropy, we force the assignments to be as close to one-hot as possible, which pushes the centroids towards their “assigned” samples. This can be seen as a form of clustering with a logistic loss. However, simply solving this problem can result in empty clusters. This is a common problem in mini-batch $k$-means (Sculley, 2010), and is usually solved by adding a reassignment phase: the centroids of the empty clusters are discarded and moved next to another non-empty centroid. In our case, we do not wish the centroids to move that abruptly, as it could disturb the training of the student. Instead, we propose to use the Sinkhorn-Knopp (SK) algorithm (Sinkhorn & Knopp, 1967), used in SwAV (Caron et al., 2020) and DINOv2 (Oquab et al., 2024). Using the SK algorithm, we obtain $a^{\prime}$, an assignment where the distribution of tokens over the clusters is near-uniform:
@@
with $\tau^{\prime}$ another temperature parameter. Note that we do not backpropagate through the SK algorithm. We adapt the loss in Problem (3), and learn the centroids $C$ by minimizing the cross-entropy between $a$ and $a^{\prime}$:
@@
We minimize this loss with AdamW (Loshchilov, 2017) alongside the training of the main model.
Positional collapse. A crucial point of self-supervised learning is avoiding trivial solutions, i.e. situations where the model minimizes the loss without learning useful features. These failure modes, also known as representation collapse, are usually addressed by adding regularization mechanisms. For example, color jittering (Chen et al., 2020b) helps prevent the model from focusing uniquely on color. While training CAPI, we observed a specific type of representation collapse as the positional encoding started outweighing the content of the patch embeddings. In the extreme case, the model learns to predict the position of the masked tokens instead of their content, resulting in a zero loss with a trivial model. In most observed cases, both content and position information are entangled in the target representation, while in the ideal scenario, the target would consist purely of semantic features.
We propose a simple solution to alleviate the problem. By running the SK separately at each position in Eq. (4), we force the joint distribution of tokens over positions and clusters to be near-uniform. Uniformity of the joint distribution directly implies zero mutual information between the clustering and the targets. This way, the modified SK ensures that our targets contain no positional information.
@@ 3.2 Predictor architecture
In this work, we use an even lighter architecture, which was initially explored in crossMAE (Fu et al., 2024) for pixel-based reconstruction. In this case, the predictor forwards only the mask tokens (Fig. 5(c)), which can access the context of the encoded patches via cross attention. Using a cross-attention predictor has two main advantages. First, it allows further efficiency gains, as we only forward a reduced set of tokens in the predictor, and we can even subsample this set of tokens. Second, mask tokens do not interact with each other in the cross-attention mechanism, i.e. each prediction is independent of other positions. This alleviates the need for multiple predictor forward passes with different prediction sets, as used in I-JEPA.
@@ 4.1 Experimental setup
For our longer experiments, we train on multiple datasets: ImageNet-1k, for comparability with previous works, ImageNet-22k, to test scaling, Places205, to test training on more diverse and less object-centric data, and finally LVD-142M, a large-scale automatically curated dataset used in previous SSL foundation models.
Model architecture. We do all our experiments with a Vision Transformer (Dosovitskiy et al., 2021) of 300M parameters (ViT-L). This architecture is widely used in various computer vision tasks, and most baselines provide a model of comparable size. We equip the vision transformer with registers (Darcet et al., 2024). These additional tokens were recently proposed as a way to add an information buffer, which enabled the model to produce smoother feature maps. For the decoder, we use 12 transformer blocks that cross-attend to the output of the encoder. This is similar to a standard transformer decoder (Vaswani et al., 2017), with the difference that we do not include self-attention layers. In this decoder, every token is forwarded independently and separately attends to the encoder output. When using a different encoder size, we align the embedding dimension, MLP ratio, and number of attention heads of the decoder to those of the encoder, and use a decoder depth equal to half that of the encoder.
@@
Baselines. We compare to the performance of previous models trained using masked image modeling: BeiT (Bao et al., 2021), MAE (He et al., 2022), data2vec 2.0 (Baevski et al., 2023), I-JEPA (Assran et al., 2023), and AIM (El-Nouby et al., 2024). To provide additional points of comparison, we report in grey the performance of iBOT (Zhou et al., 2022) and DINOv2+reg (Oquab et al., 2024; Darcet et al., 2024), who use a DINO loss to stabilize a MIM objective.
@@
Table 1: Ablation study of the main parameters and design choices in our algorithm. We report both image segmentation and classification. We highlight the default setting in gray, and bold the best-performing solution. An in-depth analysis of these results is provided in Sec. 4.2.

### v3

@@ 1 Introduction
In this work, we focus on online latent masked image modeling. We isolate it from other stabilizing objectives, and systematically study the design choices it involves. We center the discussion around three aspects of the masked image modeling principle: the target representation, the formulation of the loss, and the architecture used to perform predictions. We show that with the right implementation, a simple masked image modeling objective can lead to features competitive with the current state of the art in SSL. In brief, our method relies on a pair of vision transformers, an online model trained with masked image modeling, and a target encoder updated via an exponential moving average of the online model (fig. 1 left). The training signal for the online model comes from the patch representations of the target encoder, which is updated through an exponential moving average (EMA). The patch embeddings of the EMA encoder are converted to soft assignments on pseudo-categories using a learned online clustering, while the student receives a partially masked image as input and is trained to predict the clustering assignments of the missing patches. Using this method, we train CAPI, a 300-million parameter visual encoder whose representations approach DINOv2’s performance while significantly outperforming previous masked image models (fig. 1 right).
Figure 1: CAPI Method overview: image patches embedded by a EMA encoder are grouped into clusters. Their assignments are then used as the training signal for the online model. The loss is purely about predicting the content of missing patches and does not rely on augmentations or a contrastive loss. Evaluation scores: we evaluate frozen representations on ADE20K segmentation with a $k$-nn and linear probe and on ImageNet-1k classification with an attentive probe. We compare to MAE, data2vec 2.0, I-JEPA, and AIM. Compared to other masked image models, CAPI achieves higher performance with fewer FLOP, scaling well with model size, and approaches the scores of DINOv2+reg.
Figure 2: Detailed overview of our method with reference tensor sizes for a ViT-L/16 model. We denote in red the parts that are trained by the main loss, in purple the parts that are trained with the clustering loss, and in blue the parts that are updated by the EMA.
@@ 2 Related Work
Self-supervised representation learning. Self-supervised learning (SSL) is a pre-training paradigm in which a model is optimized to solve a pretext task on unlabeled data, often collected at scale. As the model is not trained to match specific human annotations, the benefit of this type of training is to produce a generalist model, that can be adapted to solve many different downstream tasks. Depending on the application, these tasks can be solved either by fully fine-tuning the mode, or by using the representations extracted with the frozen model. In the SSL literature, some works have focused mainly on full fine-tuning (He et al., 2022; Huang et al., 2023), while others have shown that frozen representations can reach excellent performance on a wide range of tasks, avoiding costly fine-tuning (Oquab et al., 2024) and generalizing to annotation-scarce domains where fine-tuning is not possible (Tolan et al., 2024; Xu et al., 2024). Historically, early work on self-supervised learning focused on hand-crafted pretext tasks such as predicting the rotation of an image (Gidaris et al., 2018), the relative position of patches (Doersch et al., 2015) or the color of a grayscale image (Zhang et al., 2016). Subsequent works pushed the field forward with methods based on clustering (Caron et al., 2018; Asano et al., 2020; Caron et al., 2020; 2021) and contrastive learning (Oord et al., 2018; Misra & Maaten, 2020; Chen et al., 2020b; He et al., 2020). Nowadays, the best self-supervised encoder inherits from these families (Oquab et al., 2024) and complements them with a masked image modeling objective (Zhou et al., 2022). However, training with both global and MIM objectives can prove difficult, as multiple components can interact negatively111 Zhou et al. (2022) showed that using a shared head for the DINO and iBOT losses gave better results at small scales. However, at large scale, Oquab et al. (2024) observed that this caused the iBOT loss to explode late into the training, significantly degrading performance. Untying the two heads prevented the two losses from competing directly, stabilizing the training. . In this work, we study the masked image modeling component in isolation, suggesting improvements to properly stabilize the optimization objective in the absence of a global term.
@@
Latent reconstruction. Concurrently to MAE, Zhou et al. (2022) proposed iBOT. To obtain a more semantic tokenization, iBOT used the online output of the model being trained as the reconstruction target for masked image modeling. This led to good representations, and the method was reused to obtain the current state-of-the-art (Oquab et al., 2024). However, the iBOT objective was very unstable and required an additional DINO (Caron et al., 2021) loss to stabilize the training. The idea of using online representations as targets was then reused in data2vec (Baevski et al., 2022) and I-JEPA (Assran et al., 2023; Bar et al., 2024). I-JEPA in particular proposed reusing the split encoder/decoder architecture of MAE (with the "decoder" being called "predictor" in I-JEPA) and removing the projection head of iBOT, to obtain a more stable objective in the latent space. The improvements in I-JEPA established a new tradeoff between stability and performance, but it was still both sensitive to hyperparameters ($-12$ points on IN-1k when changing the “target scale” from $[0.15,0.2]$ to $[0.125,0.2]$ (Assran et al., 2023)) and weaker than DINOv2 (81.1 on ImageNet-1k, 5.4 below DINOv2 while using a model twice bigger).
Clustering in self-supervised learning. Our approach is also related to methods that use clustering for self-supervised learning. First of this line of work, DeepCluster (Caron et al., 2018) proposed using a simple $k$-means to obtain pseudo-labels. Subsequently, SwAV (Caron et al., 2020) introduced an online clustering to replace the offline $k$-means. DINO (Caron et al., 2021) built on SWaV, and made the clustering be implicitly learned by the MLP projection head of the student. In a masked image modeling setup, iBOT (Zhou et al., 2022) proposed to reuse the DINO projection head for masked image modeling, implicitly using a clustering as target. Other recent approaches have used clustering for masked image modeling. Gidaris et al. (2024) used codebook assignments as target for a masked image modeling training, and complement it with a global loss, obtaining a formulation similar to iBOT where the clustering aspect appears more clearly. The codebook, however, is obtained by randomly sampling previous batches, whereas we attempt to learn a more optimal codebook with an online clustering. Most similar to our approach is the work of Salehi et al. (2024). They proposed a masked video modeling setup where the target encoder is a patch-wise MLP trained via gradient descent at the same time as the predicting model, via a loss comparing the target representation to the prediction of the other model. To prevent collapse, they use a Sinkhorn-Knopp algorithm to rebalance the assignments of the patches to the prototypes, obtaining a form of clustering. Both the clustering aspects and the Sinkhorn-Knopp algorithm are similar to our approach, but the target encoder is significantly different. In our study, we focus on cases where the target is the latent representation of the online model or its EMA, while they learn an entirely separate MLP to encode the targets. We also focus on images, while they focus on videos.
@@ 3 Approach
Figure 4: The target representations commonly used in MIM. (a) In the simplest case, the pixel values of the masked patches are flattened and directly used as targets. (b) An alternative is to use the output of a frozen encoder as targets, to obtain a more semantic representation. (c) To alleviate the need for a preexisting pretrained model, the target encoder can be an EMA of the online model. We focus on this case.
Overview of training. In short, our main design choices are: first, we reconstruct images in latent space with a teacher-student framework, following iBOT. As the online latent MIM framework has already shown very promising results, we restrict our focus to this specific setup. Then, we formulate our loss using a clustering component, inspired by SeLa and SwAV, and draw inspiration from the regularization methods they introduce, in particular the Sinkhorn-Knopp (Sinkhorn & Knopp, 1967) rebalancing from SeLa (Asano et al., 2020). Finally, we employ a cross-attention predictor model, separate from the encoder, to perform reconstructions, as in crossMAE (Fu et al., 2024).
Our encoder and the predictor are transformers (Dosovitskiy et al., 2021) and, during pre-training, they operate in tandem as the student (fig. 1 left). The teacher is an EMA of the encoder. The typical values for one training iteration using square images of side 224 pixels and a model with patch size 16 are:
@@ 3.1 Clustering-based loss formulation
Some self-supervised learning approaches like SwAV and DINO can be interpreted as a form of latent clustering. They employ a cross-entropy loss between the student and teacher output distributions. These distributions, produced by a linear or MLP head, can be seen as soft cluster memberships, where the centroids correspond to the prototypes. Replicating the DINO objective, iBOT proposes to pass the student predictions through an MLP head and to pass the teacher embeddings through the EMA of this head. We argue that in iBOT, in the context of MIM, the clustering interpretation of the head is not valid any more, because the loss is not done between two versions of the same representation. While in DINO both targets and predictions are [CLS] tokens, in iBOT the targets are patch tokens while the predictions are special [MSK] tokens. This can also be seen as a distribution mismatch: the MLP head of the student is trained with [MSK] inputs but is instead applied to regular patch tokens in the teacher. This mismatch would be even stronger in an asymmetric architecture as in MAE or I-JEPA, where the targets and predictions come from two different networks (see the split design in fig. 6(b)). In practice, the iBOT loss is very unstable: without the stabilizing effect of the DINO loss, the iBOT formulation is unable to bootstrap itself and results in trivial representations (see table 1(c)). Our proposition is simple: to come back to a valid clustering interpretation and, hopefully, avoid such instabilities, we explicitly use a clustering on top of the EMA network to create the target representations. This decouples the training of the teacher projection from the student’s and the training remains stable even in the absence of an additional global loss for stabilization.
(a) Direct nonparametric loss function
@@
(c) Loss after clustering assignments
Figure 5: The different loss formulations considered here. We depict in red the flow of the gradient. (a) The loss can simply be a nonparametric function, such as MSE (in MAE) or Huber loss (in I-JEPA). (b) In iBOT and DINOv2, the prediction is passed through an MLP head, and the target is passed through an EMA of this head. The loss is the cross-entropy of their outputs. (c) We propose using a clustering to obtain the targets. The loss is the cross-entropy between the predictions and the assignments.
@@
(BeiT, iBOT, SimMIM, data2vec)
(b) Split self-att. predictor
@@
Figure 6: The different predictor architectures discussed in the paper. Here, each box represents a transformer. The black lines represent the residual stream for a token. (a) In the fused architecture, the predictor and the encoder are the same transformer, which takes as input both the visible and mask tokens. (b) In the split architecture, the encoder processes only the visible tokens, while the predictor processes both the visible and the mask tokens. (c) The split architecture can be further refined by removing the visible tokens from the predictor. The predictor processes only the mask tokens, and uses cross-attention to access the context of the visible tokens.
@@
where $a_{i}^{(k)}$ is the $k$-th coordinate in $a_{i}$. By minimizing the entropy, we force the assignments to be as close to one-hot as possible, which pushes the centroids towards their “assigned” samples. This can be seen as a form of clustering with a logistic loss. However, simply solving this problem can result in empty clusters. This is a common problem in mini-batch $k$-means (Sculley, 2010), and is usually solved by adding a reassignment phase: the centroids of the empty clusters are discarded and moved next to another non-empty centroid. In our case, we do not wish the centroids to move that abruptly, as it could disturb the training of the student. Instead, we propose to use the Sinkhorn-Knopp (SK) algorithm (Sinkhorn & Knopp, 1967), used in SeLa (Asano et al., 2020) and DINOv2 (Oquab et al., 2024). Using the SK algorithm, we obtain $a^{\prime}$, an assignment where the distribution of tokens over the clusters is near-uniform:
@@
with $\tau^{\prime}$ another temperature parameter. Note that we do not backpropagate through the SK algorithm. We adapt the loss in eq. 3, and learn the centroids $C$ by minimizing the cross-entropy between $a$ and $a^{\prime}$:
@@
We minimize this loss with AdamW (Loshchilov, 2017) alongside the training of the main model. As the clustering loss and the MIM loss depend on two disjoint sets of parameters, we optimize them with two distinct optimizers without interference.
Positional collapse. A crucial point of self-supervised learning is avoiding trivial solutions, i.e. situations where the model minimizes the loss without learning useful features. These failure modes, also known as representation collapse, are usually addressed by adding regularization mechanisms. For example, color jittering (Chen et al., 2020b) helps prevent the model from focusing uniquely on color. While training CAPI, we observed a specific type of representation collapse as the positional encoding started outweighing the content of the patch embeddings. In the extreme case, the model learns to predict the position of the masked tokens instead of their content, resulting in a zero loss with a trivial model. In most observed cases, both content and position information are entangled in the target representation, while in the ideal scenario, the target would consist purely of semantic features. Although this problem does not arise in all training runs, it can be a significant issue when it does. We show in fig. 11 an example of feature maps of some early models affected by this phenomenon.
We propose a simple solution to alleviate the problem. By running the SK separately at each position in eq. 4, we force the joint distribution of tokens over positions and clusters to be near-uniform. Uniformity of the joint distribution directly implies zero mutual information between the clustering and the targets. This way, the modified SK ensures that our targets contain no positional information.
@@ 3.2 Predictor architecture
In this work, we use an even lighter predictor architecture, which was initially explored in crossMAE (Fu et al., 2024) for pixel-based reconstruction. In this case, the predictor forwards only the mask tokens (fig. 6(c)), which can access the context of the encoded patches via cross attention. Using a cross-attention predictor has two main advantages. First, it allows further efficiency gains, as we only forward a reduced set of tokens in the predictor, and we can even subsample this set of tokens. Second, mask tokens do not interact with each other in the cross-attention mechanism, i.e. each prediction is independent of other positions. This alleviates the need for multiple predictor forward passes with different prediction sets, as used in I-JEPA.
@@ 4.1 Experimental setup
For our longer experiments, we train on multiple datasets: ImageNet-1k, for comparability with previous works, ImageNet-22k, to test scaling, Places205, to test training on more diverse and less object-centric data, and finally LVD-142M, a large-scale automatically curated dataset used in previous SSL foundation models. We refer the reader to Oquab et al. (2024) for more details on the curation process.
Model architecture. We do all our experiments with a Vision Transformer (Dosovitskiy et al., 2021) of 300M parameters (ViT-L). The model has a patch size of 14, a hidden size of 1024, and 24 transformer blocks. Note that in ablation experiments, the patch size used is 16. This architecture is widely used in various computer vision tasks, and most baselines provide a model of comparable size. We equip the vision transformer with registers (Darcet et al., 2024). These additional tokens were recently proposed as a way to add an information buffer, which enabled the model to produce smoother feature maps. For the predictor, we use 12 transformer blocks that cross-attend to the output of the encoder. This is similar to a standard transformer predictor (Vaswani et al., 2017), with the difference that we do not include self-attention layers. In this predictor, every token is forwarded independently and separately attends to the encoder output. When using a different encoder size, we align the embedding dimension, MLP ratio, and number of attention heads of the predictor to those of the encoder, and use a predictor depth equal to half that of the encoder.
@@
Baselines. We compare to the performance of previous models trained using missing context prediction: BeiT (Bao et al., 2021), MAE (He et al., 2022), data2vec 2.0 (Baevski et al., 2023), I-JEPA (Assran et al., 2023), and AIM (El-Nouby et al., 2024). Although AIM is not a masked image modeling method, we believe the similarity between masked image modeling and autoregressive modeling makes it a relevant comparison. To provide additional points of comparison, we report in grey the performance of iBOT (Zhou et al., 2022), MIM-refiner (Alkin et al., 2024) (dBOT-refined) and DINOv2+reg (Oquab et al., 2024; Darcet et al., 2024), who use a DINO or contrastive loss to complement a MIM objective. When multiple model sizes are available, we report the model closest in size to our models (ViT-L/14). In the specific case of MIM-refiner, multiple models were published with different MIM pretraining bases. We select the best-performing model on our evaluations, which turns out to be dBOT-refined in all cases. Finally, we include DINO, MoCov3 and MSN (Caron et al., 2021; Chen et al., 2021; Assran et al., 2022) as representative of non-MIM SSL methods.
@@
Table 1: Ablation study of the main parameters and design choices in our algorithm. We report both image segmentation and classification. We highlight the default setting in gray, and bold the best-performing solution. We report ADE20K $k$-nn segmentation and ImageNet-1k attentive probing classification scores. An in-depth analysis of these results is provided in section 4.2.